The force of resistance on an object moving through a fluid is proportional to the square of its velocity. The initial velocity of an object of mass 1 kg moving through the fluid is 40 ms^−1. After 5 s the velocity reduces to 20 ms^−1. Calculate its velocity after
10 s.
1
Expert's answer
2019-04-30T09:45:25-0400
"ma=-kv^2"

"m\frac{dv}{dt}=-kv^2"

"\frac{dv}{v^2}=-\frac{k}{m} dt"

"\intop\frac{dv}{v^2}=-\frac{k}{m} \intop dt"

"-\frac{1}{v_0}=-\frac{k}{m} t +C"

The initial condition:


"-\frac{1}{40}=C\implies C=- 0.025"

We have:


"-\frac{1}{v}=-\frac{k}{m} t -0.025\implies \frac{1}{v}=\frac{k}{m} t+0.025"

After 5 s:


"\frac{1}{20}=\frac{k}{1} 5+0.025\implies k=0.005"

After 10 s:


"\frac{1}{v}=\frac{0.005}{1} 10+0.025"

"v=13.3\frac{m}{s}"